OpenStudy (anonymous):

solving logarithm : How do I solve this? -10 + log3 (n+3)=-10

sam (.sam.):

Where you're not sure?

OpenStudy (anonymous):

This is a problem that a friend of mine sent me, asking me for help. I know not how to solve this sort of problem.

OpenStudy (anonymous):

So I thought that I could ask in here.

sam (.sam.):

Ok, you can eliminate -10 by adding 10 both sides \[-10 + \log_3 (n+3)=-10\] \[\log_3(n+3)=0\] Then use logarithm rule \[\log_ba=c \leftarrow \rightarrow a=b^c\]

OpenStudy (anonymous):

Thank you. :)

sam (.sam.):

Welcome

OpenStudy (anonymous):

This is what he sent me. logba=c←→a=bc ?

OpenStudy (anonymous):

So which one is ba and which one is c?

sam (.sam.):

You can directly see it a=n+3 b=3 c=0

sam (.sam.):

Using \(\log_ba=c \leftarrow \rightarrow a=b^c\) \[n+3=3^0 \\ \\ n+3=1 \\ \\ n=?\]

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

Thank you again.